### v4

Abstract
Security issuance through intermediaries is subject to informational and agency-related frictions. However, separating their effects on securities has been difficult. We estimate those effects separately using SPAC data. To that end, we identify “premium” investors who produce value-relevant information. Their participation is associated with lower liquidation risk and higher returns. In contrast, “non-premium” investors engage only in quid pro quo arrangements. They receive high returns from an intermediary (quid) for participating in weaker future deals initiated by that intermediary (quo). Thus, quid pro quo is not pure agency cost; it includes transfers enabling more firms to go public.
Keywords: Security Issuance, Institutional Investors, SPAC, Underpricing JEL classification: G14, G32, G34.
1 Introduction
Firms issuing securities through intermediaries face two frictions. First, an informational gap between them (issuers) and investors may prevent securities from being sold at full value. Second, there may be an agency problem if intermediaries do not maximize sale proceeds but instead engage in quid pro quo arrangements with investors. Specifically, intermediaries may favor certain buyers (investors) at the expense of the issuers, expecting their favoritism to be returned through profitable future transactions.
Ritter and Welch (2002) point out that both frictions coexist and emphasize the need and difficulty in separating those frictions. In a survey paper, Lowry et al. (2017) similarly pose this separation as an unresolved question, using IPO underpricing as an example:
“Does [IPO underpricing] reflect a reward to institutions for sharing value-relevant information with underwriters, or does it reflect a quid pro quo arrangement where the bank is rewarding its best clients? [..] This is an immensely important issue and one that we hope future research will be able to address.”
We revisit this problem and separately estimate the effects of these frictions in the context of special purpose acquisition companies (SPACs)–publicly traded shell companies formed by sponsors to acquire or merge with private targets to take them public.
Like underwriters in a traditional IPO, sponsors are intermediaries between a target firm wanting to go public and investors. In both cases, an intermediary negotiates with a private company to determine the offer price, number of shares sold, and allocation across investors. Likewise, they profit if and only if the merger is successful.
In our setting, we distinguish between two types of investors: “premium” and “non-premium” (defined shortly below). We find evidence consistent with premium investors acting in the information production role. Their certification is associated with successful mergers and higher announcement returns. Non-premium investors engage only in quid pro quo arrangements. Specifically, when non-premium investors are allocated shares from a “hot” deal, they participate in a weaker deal initiated by the same sponsor in the future. Thus, interestingly, this quid pro quo is not all value-destroying. It partly enables a wider range of issuers to access public markets.
We can isolate the effects of informational frictions separately from the agency friction because the SPAC setting offers several research design advantages over the traditional IPO. First, a SPAC divides the process of going public into distinct phases. In the initial stage, the sponsor raises capital from investors. A SPAC starts as a newly formed shell company without an operating business, so there is limited scope for information asymmetry (about underlying business) compared to a traditional IPO that starts with an operating business.111Consistent with this, in our sample, we find a small (1.3%) underpricing at SPAC IPO. The next phase consists of a “business combination” in which the SPAC merges with a private business and introduces new institutional investors known as “PIPE” (private investment in public equity), which affords us the ability to isolate the effects of PIPE investors.
Second, legally, SPACs are publicly traded firms. Consequently, we observe the allocation to each PIPE investor and the effect of major events (e.g., business combination) on share prices. These data enable us to measure the return history for every sponsor-PIPE pair and create an in-sample measure of past relationships.222In traditional IPOs, allocations to institutional investors are unreported, and the quarterly 13F filings are insufficient to quantify favoritism if preferred clients “flip” shares in the secondary market. Furthermore, 13F filings do not distinguish shares allocated at the IPO from shares purchased in the secondary market. Third, the quid pro quo channel is easier to observe in SPACs than traditional IPOs. Unlike the intermediaries (investment banks) in traditional IPOs, SPAC sponsors typically do not operate in other lines of business (e.g., trading and wealth management). To the extent that sponsors earn quid pro quo from investors in future SPAC deals, we can observe that in our data.
Using data on all 1,072 SPACs from January 2010 to December 2021, we classify PIPE investors into premium and non-premium types based on participation frequency and investment size using k-means clustering. All else equal, frequent and large investors are more likely to produce value-relevant information than others (Benveniste and Spindt, 1989).
A challenge in estimating the effects of premium investors is that the set of investors working with a sponsor is non-random. Better or more experienced sponsors may choose to work with better premium investors. If so, the estimated effects of premium investors could be biased because it would include the effect of assortative matching. So, we estimate fixed effects for each sponsor and investor following the methodology of Abowd, Kramarz and Margolis (1999) and use those fixed effects in our regression analysis. Our identifying assumption is that conditional on the fixed effects, the sponsor and premium investor matching is random.
First, we find evidence consistent with premium investors offering value-relevant information and playing certification roles. Their participation is associated with significantly less share redemption by public investors and a large positive announcement effect on the stock price. Our estimates suggest that one standard deviation (32.85 pps) increase in premium investors’ participation is associated with a 5.91 pps decrease in the redemption rate (10.5% of the sample mean) and a 4.40 pps increase in the announcement-day return (117.7% of the sample mean).
Next, we consider the agency friction. We note that sponsors profit an average of $126.9 million when the merger is successful; otherwise, they lose their initial investment, which averages $8.2 million. Therefore, the sponsors benefit from investors’ participation, which improves the chances of a successful merger. In turn, sponsors may have incentives to offer discounted shares to those investors in strong deals in exchange for an implicit promise to help them out with weaker deals in the future. We find evidence that only the non-premium investors engage in quid pro quo arrangements.
In particular, we use the instrumental variables approach to document three ways quid pro quo manifests in the data. First, for each sponsor and non-premium investor pair, money left on the table (acronym, “MLOT”) exhibits mean-reversion, i.e., non-premium investors profit less in the current deal when they have high returns on past deals, and vice versa. In particular, one dollar above average MLOT in the last deal leads to a $2.72 below average MLOT from the current deal. We interpret this mean-reversion as the sponsor building a relationship with some non-premium investors and drawing it down when needed.
Second, when the deal is weak or “cold,” i.e., when the offer share price at IPO is higher than the price on the closing day of the merger, the sponsor allocates shares disproportionately to those who earned strong returns from that sponsor in the past. Moreover, the sponsor allocates non-premium investors with a strong relationship more securities when the deals are weak than when they are strong.
Third, a sponsor with strong relationships with non-premium investors is more likely to succeed. More precisely, if sponsors allocate one standard deviation more MLOT above average ($300 million) to non-premium investors, the liquidation probability of the current deal falls by 48.93 pps. As a reference, the average liquidation rate in our sample is 17.82%, so this decrease is economically meaningful. We also implement several robustness exercises, including placebo exercises, and find that our results are consistent with our hypotheses.
Thus, we establish a new institutional feature in security issuance based on the relationship between the intermediary and institutional investors. In particular, we provide evidence that the quid pro quo arrangement enables cross-deal subsidies from high-quality target firms that will allow low-quality firms to go public. In contrast, the existing literature treats quid pro quo as a pure agency cost borne by issuers.
In summary, premium and non-premium investors play complementary roles and are paid accordingly. A premium investor, on average, is paid $6.96 million to produce information for the current target. A non-premium investor is paid $3.17 million to reduce liquidation risk in the future when the deal is weak, and premium investors are less likely to participate.
We contribute to a long line of research that uses underpricing to infer the importance of agency and informational problems. Underpricing has been argued to be an implication of information asymmetry between investors (Rock, 1986) or between investors and intermediaries (Benveniste and Spindt, 1989). In contrast, Beatty and Ritter (1986); Loughran and Ritter (2002) and Loughran and Ritter (2004) argue that underpricing is pure unproductive rent. One debate has been over the nature of ‘partial adjustment,’ i.e., underwriters partially adjust to good information revealed by investors but fully adjust to bad information. Hanley (1993) presents supporting empirical evidence. Also see Bradley and Jordan (2002) and İnce (2014) for more. We complement this literature by examining broader outcomes and showing that information and quid pro quo arrangements are present.
Our findings about the differential effects of premium and non-premium investors complement the findings in Wang and Yung (2011) and Chemmanur, Hu and Huang (2010). Additionally, our finding that information production coexists with quid pro quo is reminiscent of Jenkinson, Jones and Suntheim (2018). Using data on the initial allocations of investment banks participating in traditional IPOs, they find evidence that investors are rewarded more when they submit price-limited bids and participate more actively in the meetings before or during bookbuilding. Both behaviors are consistent with an informational role. However, they also find that allocations are affected by how much revenue the investor generates for the bank, which is consistent with the quid pro quo argument. Our notion of quid pro quo differs slightly from the literature. Jenkinson, Jones and Suntheim (2018) document that investors build a relationship with a bank to get underpriced shares. We instead document that sponsors build relationships with non-premium PIPE investors to use them to complete future deals.
We also contribute to a nascent literature that seeks to understand the role and functioning of SPACs. For instance, Gahng, Ritter and Zhang (Forthcoming) study the cost and benefit of using SPACs to go public; Gofman and Yao (2022) study the conflict of interest and associated harm from directors serving on the boards of multiple SPACs; and Klausner, Ohlrogge and Ruan (2022) analyze the structure of SPACs and the associated costs.
2 A Primer on SPACs
A SPAC is formed when a sponsor begins the SEC-mandated initial public offering process. It is, first and foremost, an IPO and must comply with all SEC IPO regulations. In addition, SPACs are subject to additional requirements as “blank check companies,” including how they use their escrow fund (described below). The SPAC aims to identify and acquire a private company to take it public.
During their IPOs, SPACs offer securities to public investors called units, typically priced at $10 per unit.333One exception is Periphas Capital Partnering Corporation, which priced its unit at $25. A unit comprises one Class A common stock share and a proportion (e.g., 1, 1/2, 1/3) of one redeemable warrant. Initially, these units trade as a whole. However, usually around the 52 calendar day following the IPO, owners have the option to either continue to hold units as a whole or trade shares and warrants separately.444After the separation, SPACs do not issue fractional warrants and only whole warrants trade. In other words, the number of warrants issued after separation equals the fraction rounded down to the nearest whole number. For example, if a unit contains one share and 1/3 of a warrant, then an investor with 1 unit receives no warrant, but an investor with 30 units receives ten warrants. Once the merger deal between the SPAC and a target is closed, a warrant gives an investor the right to purchase additional shares of common stock. The strike price is typically set well above the original SPAC price; a common choice is $11.50 per share. Any shares created in this manner automatically convert into ordinary shares after a merger.
Sponsors purchase shares before SPACs go public. These shares are called the “founder shares” or the sponsor’s “promote”. Unlike Class A common shares in the units owned by public investors, the sponsor’s shares become worthless if the SPAC liquidates.
A SPAC generally has 24 months from its IPO to identify a target company, negotiate a deal, and complete the merger. Failure to complete the merger on time leads to its liquidation, and the SPAC must redeem all public shares and pay its investors with the funds held in the trust account, including any interest earned.
Once the SPAC locates a target, it uses proceeds from the trust fund for the merger. Until then, the trust fund can be used only to pay public shareholders who redeem their shares. The sponsor creates a “risk capital” to meet operating expenses by selling private placement warrants. These warrants are irredeemable when the SPAC liquidates and are subject to transfer restrictions.
Next, SPAC shareholders vote on the deal, and the transaction occurs only if a majority votes in favor. Shareholders can also redeem their shares, typically having at least 20 business days before the merger is complete. The merger agreement with the target company requires that the SPAC have a minimum net worth or a certain amount of cash (per the negotiation) as a closing condition. Sometimes, the funds raised in the SPAC IPO are sufficient to purchase the target the sponsor has located. More typically, the sponsor needs to raise additional funds. Because the SPAC is a publicly traded firm, these funds are called PIPEs.
PIPE investors are accredited institutional investors, e.g., hedge funds and mutual funds, that agree to purchase securities issued by a company. Such an agreement requires the company to file a resale registration statement (in the future) to allow investors to resell their securities. Compared to raising capital in the public market, PIPE investments tend to have lower transaction costs, and companies can disclose transaction details after receiving commitments from PIPE investors.
We categorize PIPE transactions based on the timing of their commitment: before, at, and after the business combination. In particular, the sponsor can invite PIPE investors before the SPAC IPO to commit to making certain investments in connection with the business combination. The PIPE transactions formed at the time of the IPO but before a target is identified are called forward purchase agreements; henceforth, they are PIPE (pre). The second type of PIPE transaction is announced along with the news about the business combination; henceforth, PIPE (at). The third type of PIPE transaction occurs after the target is announced but before the merger is completed; henceforth, PIPE (post).
A SPAC can have any combination of these three types of PIPEs. For example, Table 1 shows an example of SPAC, CF Capital Acquisition Corp V, which had all three PIPEs.
| Key Events | Dates |
| S-1 Filed and PIPE (pre) Announced | 08-Jan-2021 |
| IPO | 29-Jan-2021 |
| Deal Announced, PIPE (at) Announced | 06-Jul-2021 |
| Proxy Filed | 12-Aug-2021 |
| PIPE (post) Announced | 18-Jan-2021 |
| Shareholder Vote | 24-Jan-2022 |
| Deal Closed | 25-Jan-2022 |
| Liquidation Deadline | 02-Feb-2023 |
Note: Key events and the dates for the SPAC CF Acquisition Corp V.
In Figure 1, we compare the timeline of a SPAC and a traditional IPO. The date most analogous to the IPO date in a traditional initial public offering is the business combination closing: the point at which a private operating firm becomes public.
Note: This figure illustrates the timelines of a traditional IPO (top) and a SPAC (bottom). The acronym “BC" stands for a business combination.
3 Testable Hypotheses
We develop testable hypotheses based on two competing paradigms – the information production and quid pro quo arrangements – that we assess in our data.
Information Production. Under the first paradigm, investors have or produce private information about the issuer and earn rents accordingly. This paradigm views a central goal of the going public process as extracting value-relevant information about the issuer and incorporating this information into the price.
Going public via SPACs involves investors in three stages (see Figure 1), and information production may have different implications at each step.
In particular, information production predicts zero underpricing during the SPAC IPO because a target has yet to be identified, and thus, there is no due diligence to perform. There is virtually no scope for investors to have private information; they exchange cash for a proportion of a trust fund of a fixed value.
In the business combination or de-SPAC phase, investors that participate in the target announcement (PIPE (at)) are the analogs to the buyers in the IPO bookbuilding models. Specifically, the sponsor finds a target and approaches PIPE (at) investors to solicit their interest. Then, the sponsor builds a book of PIPE demand, sets prices, and allocates shares.
However, these investors may differ in their abilities to produce information. For instance, Benveniste and Spindt (1989) suggest that premium investors who frequently access the market are likely more informed than others. In our context, such investors have private information about the quality of the issuing (target) firm and overall conditions affecting the market valuation. Conversely, non-premium investors are less informed and consequently have less impact on prices. Under these assumptions, the securities of the target firm may be underpriced to compensate premium investors for truthfully disclosing their private information about the target.
Furthermore, a premium PIPE (at) investor is incentivized to purchase more shares when its private information indicates the de-SPAC deal is profitable. Additionally, sponsors prioritize allocating to premium investors to encourage them to disclose their information. Consequently, a larger allocation to premium investors signals the good quality of the deal, which reduces redemption by public investors and leads to a positive market reaction. However, as non-premium PIPEs have no certification role, their participation should neither affect redemption nor lead to higher announcement-day returns.
Conversely, we posit that PIPE (pre) and PIPE (post) investors do not play certification roles. In the case of PIPE (pre), there has yet to be any information about the target, so there is nothing to certify. PIPE (post) investors are used less often than PIPE (at) investors and can be used to fill the financing gap caused by redemption. If the sponsor intended a PIPE investor to be used for certification purposes (as opposed to “gap filling”), the sponsor would have announced the participation of the PIPE contemporaneously when the deal was announced rather than waiting until after public investors reacted.
Suppose a premium PIPE (at) investor acts purely as an information producer and is under no obligation to the sponsor to maintain a relationship. In this case, the PIPE participates in deals it likes and is absent otherwise. Such an arrangement should not generate any mean reversion when the unit of observation is sponsor-PIPE.
Hypothesis 1.
Under the information production paradigm, the following should hold:
1.
Allocations to premium PIPE (at) investors are associated with lower redemption and positive announcement-day returns.
2.
For a sponsor-premium PIPE (at) pair, MLOT is not mean-reverting.
Quid Pro Quo. We now discuss the implications of the quid pro quo paradigm. Under this paradigm, the seller of securities acts through a self-interested intermediary. Rather than selling securities at full price to premium investors, the intermediary gives discounts to some buyers, with an implicit understanding that this favor is rewarded through profits on future SPAC transactions.
The SPAC IPO presents the underwriter with such an opportunity. Underwriters likely deal with institutional investors more frequently (and in more varied capacities) than sponsors. As an intermediary, the underwriter may favor institutional investors (buyers) over sponsors (sellers), just as in a traditional IPO, resulting in SPAC underpricing.
The business combination phase presents another opportunity to observe favoritism, except the actors are swapped. The issuer is the private firm rather than the SPAC because the private firm, not the SPAC, issues new securities. The buyers are the PIPE investors who infuse new capital in exchange for securities. The intermediary is the sponsor, which negotiates terms and profits with PIPE investors only if the deal is completed.
Suppose that the underpricing at the business combination phase is not because of information production but because the sponsor allocates rents to PIPE (at) investors. Then, the sponsor should receive returns from these investors in the future. Then, including PIPE investors in a SPAC would not benefit the SPAC. Therefore, allocation to any PIPE (at) investors should not be associated with either redemption rates or announcement-day returns.
In any given deal, sponsors can favor particular PIPE investors and insist on favorable terms from the target firm. Favoritism may manifest differently at this stage than in traditional IPOs, where underwriters operate in many business lines, including trading and wealth management. An institutional investor receiving underpriced shares in a traditional IPO may repay the underwriter by providing business through other channels. In contrast, the sponsors operate only in the SPAC industry, and the primary objective of a SPAC sponsor is to close the existing fund and start future SPACs
We hypothesize that quid pro quo arrangements are employed only with non-premium PIPE investors. A sponsor showing favoritism to a particular non-premium PIPE investor expects reciprocity in future deals. Specifically, non-premium PIPE investors only participate in an underpriced deal (acting as a tax on the current private firm) in exchange for implied participation in a future weak deal. If a PIPE is involved only in “strong” deals, then the relationship is one-sided and involves “quid” with no “quo.”
While a long-term relationship (e.g., Levin, 2003; Halac, 2012) between a non-premium PIPE investor and a sponsor can provide necessary incentives to the investors to participate in weaker future deals, the observed interactions in any data are likely finite. Then, the question remains whether non-premium investors have incentives to participate in weaker future deals. In our context, it is appropriate to view non-premium investors as long-lived players who interact with sponsors who are short-lived players with perfect public monitoring. Such investors do not need to have repeated interactions with a particular sponsor to have the incentive to honor their commitment to participate in weak future deals (Fudenberg and Levine, 1994). If they renege to join in a weak deal led by a sponsor from whom they have received favorable returns, they risk losing future SPAC business.
Furthermore, under quid pro quo arrangements, the sponsor benefits from creating goodwill among investors, hoping to “draw down” on this goodwill in the future when the deals are weak. So, we hypothesize that a SPAC’s liquidation risk is negatively correlated with the amount of MLOT investors have received in the past from the sponsor of this SPAC.
Hypothesis 2.
Under the quid pro quo paradigm, the following should hold:
1.
Allocations to non-premium PIPE (at) investors are uncorrelated with redemption and announcement-day returns.
2.
For a given sponsor -non-premium PIPE (at) pair, MLOT is mean-reverting.
3.
When a deal is weak, the sponsor is likelier to allocate shares to non-premium PIPE investors who earned large past MLOT.
4.
SPACs are less likely to liquidate when the sponsor has earned goodwill through high past MLOT. This connection occurs through only non-premium and not premium PIPEs.
4 Data
We consider all 1,072 SPACs traded in major exchanges between January 2010 and December 2021. See Figure A4 for the monthly scatter plot in Figure A4, which includes the so-called boom and bust cycle for SPACs. We construct our sample by merging five datasets: (i) the SPAC Research,555https://www.spacresearch.com/. (ii) the PrivateRaise–SPACs Platform,666https://www.privateraise.com/about/platform_spac.php. (iii) the PrivateRaise–PIPEs Platform (iv) Bloomberg data on SPAC prices, and (v) Super 8-K filings from SEC.777https://www.sec.gov/edgar/search-and-access.888Thus, we exclude SPACs traded in the over-the-counter markets. After the merger between a SPAC and a target is complete, the SPAC is delisted. The new (merged) company may be assigned a new ticker symbol. We use Bloomberg-assigned unique identifiers to match across samples.
The SPAC Research database and PrivateRaise’s SPACs platform database contain information on SPACs, including important event dates. PrivateRaise’s PIPE platform database includes comprehensive information on PIPE transactions. The platform covers more than 20,000 placement profiles of PIPE transactions. It also provides contract terms, investment amounts, types of transactions, prices, identities, and allocations of the PIPE investors.
For these 1,072 SPACs, we manually merged the first two databases. The issuer’s name recorded in the PIPE transactions can be either that of SPAC or the post-merger company. In those cases, we use both names to match the SPAC sample with the PrivateRaise database. We also use the total PIPE amount from PrivateRaise to cross-check the match to ensure match accuracy. If there are inconsistencies, we use the SPAC’s SEC filings documents to record the correct amount. We then merge our sample with Bloomberg’s pricing data; warrant prices are unavailable in CRSP. Our matching process yields detailed information on transactions for all three types of PIPE investments for each SPAC in our sample.
In summary, we observe each SPAC’s life-cycle status, IPO structure, IPO outcomes, sponsor and management team, redemption, merger details, merger outcome, PIPE transactions, and price movement across events. As of February 17, 2023, 473 of these 1,072 SPACs had successfully merged, 128 had announced that they had identified a target, 280 were in the process of identifying targets, and 191 had failed and liquidated.
The number of SPAC shares changes over time on account of i) investors exercising warrants, ii) redemption of shares, iii) securities purchased by PIPES, and iv) the fact that founders sometimes voluntarily forfeit securities. Consequently, the number of shares listed in the IPO prospectus is no longer current as of the business combination date. We use Super 8-K (Super 20-F for foreign issuer) filings to collect information on the number and types of outstanding securities on the first listing day of the post-merger company. Information about securities at the time of the SPAC IPO may not be comprehensive either because the sponsors have forfeited their founder shares and private placement securities or the units that contain rights would have become common shares or both. We manually collect information from the Super 8-K (or Super 20-F) filings to accurately measure the number of outstanding securities post-merger. SEC mandates that SPACs file Super 8-K within four business days after the merger is complete, and it records the shares and warrants held by sponsors and public investors.
Table 2 provides summary data. On average, SPACs raise $272.42 million proceeds during their IPO, with a median of $240 million and 90th percentile at $440 million. A typical SPAC unit contains 0.44 warrants and 0.01 rights. During the IPO process, SPACs’ underwriters choose, on average, to exercise the 11.59% over-allotment option to meet the additional demand.
SPACs take, on average, 361 days to find a target and an additional 168 days to complete the merger process. After targets’ announcements, 56.54% of SPAC’s public shareholders redeem their shares on average. Table 2 also displays information about the participation of PIPE investors for 473 SPACs that completed the merger process during our sample period. At the time of their IPOs, 10% of SPACs have pre-committed investment from PIPE investors, i.e., PIPE (pre). When SPACs announce their targets, 66% of them also announce investments from PIPE (at), and 12% of SPACs obtain additional investments from PIPE (post) after the target announcement.
Returns. As seen in Table 2, the average return for SPAC IPO investors on the first trading day is 1.3%, and when the target is announced, SPAC investors receive an average return of 3.74%.
Money Left on the Table. We use money left on the table (“MLOT”) to measure the profit accrued by PIPE investors. The MLOT denotes the difference between the closing-day price on the first trading day of the post-merger company and the offer price presented to PIPE investors by the SPAC, and this difference is multiplied by the number of securities held by the PIPE investors. This calculation considers both common shares and warrants acquired by the PIPE investors. To determine the rate of underpricing, we normalize the MLOT by PIPE investors’ total investment amount.999Although the contracts with PIPE investors are signed before the merger is closed, they represent a pre-commitment price, implying that the PIPE investors deploy their capital only at the merger’s closure. This arrangement mirrors a ”one-day” return that traditional IPO investors experience. We carry out the MLOT and underpricing calculations for all three forms of PIPE investors (pre, at, and post).
| Mean | Std.Dev. | P10 | P50 | P90 | Obs. | |
| SPAC Measures | ||||||
| IPO proceeds (mm USD) | 272.42 | 218.22 | 100.00 | 230.00 | 440.00 | 1,072 |
| No. of Warrant | 0.44 | 0.24 | 0.20 | 0.50 | 0.75 | 1,072 |
| No. of Right | 0.01 | 0.03 | 0.00 | 0.00 | 0.00 | 1,072 |
| Overallotment (%) | 11.59 | 5.37 | 0.00 | 15.00 | 15.00 | 1,072 |
| Listing-day return (%) | 1.30 | 3.02 | -0.70 | 0.40 | 4.00 | 857 |
| Days searching | 361.71 | 215.79 | 114.00 | 329.50 | 677.00 | 614 |
| Days remaining | 340.98 | 189.87 | 120.00 | 303.00 | 611.00 | 614 |
| Announcement-day return (%) | 3.74 | 12.56 | -0.68 | 0.34 | 9.95 | 586 |
| Redeemption(%) | 56.54 | 37.49 | 0.00 | 70.00 | 97.00 | 445 |
| Days deSPAC | 167.99 | 67.06 | 100.00 | 160.00 | 241.00 | 474 |
| PIPE Participation | ||||||
| {PIPE (pre)} | 0.10 | 0.30 | 0.00 | 0.00 | 0.00 | 473 |
| {PIPE (at)} | 0.66 | 0.47 | 0.00 | 1.00 | 1.00 | 473 |
| {PIPE (post)} | 0.12 | 0.33 | 0.00 | 0.00 | 1.00 | 473 |
| {Premium PIPE (at)} | 0.50 | 0.50 | 0.00 | 1.00 | 1.00 | 473 |
| {Non-Premium PIPE (at)} | 0.65 | 0.48 | 0.00 | 1.00 | 1.00 | 473 |
| No. of PIPE (at) | 10.76 | 11.83 | 0.00 | 8.00 | 26.00 | 472 |
| No. of Premium PIPE (at) | 3.02 | 4.36 | 0.00 | 1.00 | 10.00 | 472 |
| No. of Non-Premium PIPE (at) | 7.75 | 8.64 | 0.00 | 6.00 | 18.00 | 472 |
| PIPE Investment/IPO Proceeds | ||||||
| PIPE (pre) (%) | 2.93 | 11.22 | 0.00 | 0.00 | 0.00 | 473 |
| Premium PIPE (at) (%) | 17.84 | 32.85 | 0.00 | 0.00 | 52.39 | 473 |
| Non-Premium PIPE (at) (%) | 41.94 | 58.64 | 0.00 | 28.57 | 101.16 | 473 |
| PIPE (post) (%) | 7.45 | 30.39 | 0.00 | 0.00 | 22.68 | 473 |
| MLOT (mm USD) | ||||||
| Public Investors | 137.35 | 323.71 | -2.96 | 33.85 | 364.26 | 162 |
| Sponsors | 109.76 | 161.45 | 23.00 | 77.30 | 210.74 | 160 |
| PIPE | 201.22 | 1,200.54 | -25.60 | 3.75 | 248.17 | 358 |
| PIPE (pre) | 29.37 | 120.76 | -10.52 | 3.17 | 35.70 | 45 |
| PIPE (at) | 193.07 | 1,135.58 | -25.36 | 4.77 | 248.17 | 303 |
| PIPE (post) | 8.07 | 42.62 | -15.74 | 0.25 | 31.76 | 56 |
| Premium PIPE (at) | 76.63 | 374.15 | -10.22 | 1.68 | 129.46 | 231 |
| Non-Premium PIPE (at) | 134.65 | 838.67 | -17.73 | 3.40 | 138.48 | 303 |
| MLOT/Investment (%) | ||||||
| PIPE | 91.26 | 472.71 | -16.50 | 4.47 | 85.80 | 358 |
| PIPE (pre) | 70.33 | 289.37 | -8.05 | 4.36 | 91.80 | 45 |
| PIPE (at) | 63.74 | 307.47 | -16.00 | 4.20 | 80.90 | 303 |
| PIPE (post) | 70.74 | 285.16 | -36.02 | 0.80 | 110.60 | 56 |
| Premium PIPE (at) | 64.05 | 291.75 | -13.00 | 5.30 | 94.90 | 225 |
| Non-Premium PIPE (at) | 64.10 | 308.46 | -16.00 | 4.20 | 80.90 | 301 |
Note: The sample includes 1,072 SPACs that went through the IPO process between January 2010 and December 2021. All variables are defined in Table B1. We exclude a PIPE (post) investor named Quinpario Acquisition Corporation 2, with unusually large returns.
Table 2 highlights the MLOT for all three types of PIPE investors. Investing in a SPAC generates an average of $201.22 million for PIPE investors. Out of the 45 SPAC deals in which they participate, PIPE (pre) investors make $29.37 million on average. From the 303 SPAC deals in which they take part, PIPE (at) investors earn $193.07 million per deal. PIPE (post) investors participate in 56 deals and earn $8.07 million per deal on average.
A SPAC deal generates an average return of 91.26%, where PIPE (pre) investors earn 70.33%, PIPE (at) investors earn 63.74%, and PIPE (post) investors earn 70.74%.
Premium PIPE Investors. In total, there are 2,005 PIPE investors in our sample. Out of them, 76 invested in PIPE (pre) transaction, 249 in PIPE (post), and 1,826 invested in at least one PIPE (at) transaction, but only a few large institutional investors participated frequently and invested significantly more than others. To capture this distinction, we use k-means clustering on the total investments in all SPACs and the number of SPACs to classify PIPE investors into two groups. We refer to a subset of PIPE investors who frequently participate as “premium” PIPE investors and the rest as non-premium PIPE investors. There are 34 premium PIPE investors in our sample (Table 3).101010As a way to corroborate this classification, we counted the number of articles written about each of these 34 premium investors in The New York Times and found that these premium investors have the highest counts. For example, excluding research universities, e.g., Columbia University, and non-financial companies, e.g., General Motors and Microsoft Inc., top ten investors include Morgan Stanley, Fidelity, USB, Wells Fargo, Blackstone, Citadel, Bain Capital, BlackRock, T. Rowe Price, and SoftBank.
| Investor | #SPAC | Volume (million USD) |
| Fidelity Management & Research Company | 69 | 4,342 |
| BlackRock, Inc. | 57 | 2,317 |
| Capital Research and Management Company | 29 | 2,231 |
| Alyeska Investment Group L.P. | 106 | 1,704 |
| Morgan Stanley | 14 | 1,238 |
| Millennium Management, LLC | 77 | 1,059 |
| Luxor Capital Group | 52 | 972 |
| Blackstone Inc. (f/k/a The Blackstone Group Inc.) | 34 | 965 |
| Koch Industries Inc. | 21 | 926 |
| Third Point, LLC | 12 | 910 |
| Hedosophia Services Limited | 23 | 910 |
| T. Rowe Price Group, Inc. | 27 | 899 |
| SoftBank Group Corp. | 16 | 795 |
| Franklin Resources, Inc. (d/b/a Franklin Templeton Investments) | 24 | 741 |
| Senator Investment Group LP | 36 | 713 |
| Moore Capital Management, Inc. | 66 | 629 |
| Neuberger Berman Group LLC | 20 | 626 |
| Baron Capital Group, Inc. | 25 | 621 |
| Citadel LLC | 51 | 604 |
| Hudson Bay Capital Management LP | 56 | 577 |
| Suvretta Capital Management, LLC | 31 | 576 |
| UBS AG | 57 | 494 |
| Wellington Management Company, LLP | 26 | 437 |
| Heights Capital Management, Inc. | 65 | 431 |
| Magnetar Capital LLC | 39 | 424 |
| D.E. Shaw & Co., L.P. | 38 | 411 |
| Monashee Investment Management LLC | 61 | 408 |
| Ghisallo Capital Management LLC | 62 | 389 |
| Jane Street Group, LLC | 49 | 368 |
| Park West Asset Management, LLC | 27 | 342 |
| Kepos Capital LP | 51 | 286 |
| Linden Advisors LP | 43 | 282 |
| Schonfeld Strategic Advisors LLC | 32 | 224 |
| BlueCrest Capital Management Ltd. | 31 | 181 |
Note: This table lists the name of premium PIPE investors identified using k-means clustering based on the number of SPAC deals and total investment volumes using data on SPACs that went through an IPO from January 2010 to December 2021.
Table 2 offers summary statistics for premium and non-premium PIPE investors. A standard SPAC deal usually involves about ten PIPE (at) investors, typically divided into three premium and seven non-premium investors. In the 231 SPAC transactions in which they participate, premium investors realize an average MLOT of $76.63 million with an average return of 64.05%.
In contrast, non-premium investors participate in 303 SPAC deals, generating an average MLOT of $134.65 million and achieving an average return of 64.10%. The outcomes for non-premium investors display greater variability in terms of dollar values.
It is important to note that within each SPAC, both premium and non-premium investors are given the same pricing terms, indicating that the difference in returns within each SPAC arises purely from the variance in their investment amounts.
This comparison is consistent with the quid pro quo arrangement where non-premium investors attain high returns on successful deals, offset by low returns on less successful ones. In contrast, premium investors see more steady returns due to their superior information. The difference between the MLOTs of premium and non-premium investors is further analyzed in Section 6.
5 Empirical Results
In this section, we investigate whether the data are consistent with the hypotheses stated in Section 3. We begin measuring underpricing at the SPAC IPO level and compare it with the traditional IPO underpricing. Then, we show that the premium PIPE investors who participate in business combinations generate value-relevant information, whereas non-premium investors engage in a quid pro quo relationship. We end by quantifying the benefit of the quid pro quo relationship in terms of lowering the liquidation risk of future SPACs.
5.1 Underpricing
5.1.1 Underpricing during SPAC IPO
We first examine underpricing during the SPAC IPO. Table 2 reports that the average equal-weighted first-day return of SPACs is 1.3%. Although our data start from 2010, most SPAC IPOs occurred between 2020 and 2021. In 2020, 248 SPACs went public, with an average equal-weighted first-day return of 1.3%; in 2021, 613 SPACs went public, and their average equal-weighted first-day return was 1.51%. Although these measures are statistically different from zero (p-values in both cases), they are small.111111The average equal-weighted first-day return for traditional IPOs was 41.6% in 2020 and 32% in 2021; see https://site.warrington.ufl.edu/ritter/files/IPOs-Underpricing.pdf (accessed 09/22/2022). In addition, SPACs almost always trade at $10 per share until the target announcement date. This uniformity is inconsistent with any information about sponsor quality revealed during the SPAC IPO.
5.1.2 Underpricing during de-SPAC
This section focuses on the underpricing associated with the de-SPAC stage. We say that a private company’s shares are underpriced if the first-day trading value of the post-merger company’s securities held by all SPAC investors (including sponsors, PIPEs, and original SPAC investors who did not redeem) exceeds the proceeds invested in the private firm.
There are at least two sources of asymmetric information at this stage. First, the target company may be better informed about its business prospects than anyone else (Smith Jr, 1986; Booth and Smith II, 1986). Second, the PIPE investors may also have better information about factors affecting the target firm than the public investors. In response to PIPE investors’ information, sponsors may underprice the securities issued to those investors as an incentive payment for their value-relevant information. We focus on the second type of information.
From Table 2, we see that, on average, a private company that uses the SPAC route to access the public market leaves $126.93 million MLOT to the sponsors and $201.22 million to the PIPE investors. Furthermore, we also see that PIPE investors who participate in different stages of SPAC are compensated differently. During the SPAC IPO, PIPE (pre) investors earn an average of $29.37 million through forward-purchase agreements, but PIPE (at) investors get, on average, a total of $193.07 million per deal. After the target is disclosed, PIPE (post) investors who participate receive an additional $8.07 million MLOT. Even among PIPE (at) investors, approximately one-third of the MLOT accrues to 34 premium investors. This difference suggests that the two types of investors have different business models, which we return to later in Section 6.
Overall, the SPAC shares offered to PIPE (at) investors, especially the premium PIPE investors, during the de-SPAC phase are underpriced. These results are consistent with Hypothesis 1 that PIPE (at) investors participating in the de-SPAC stage are privately informed about the quality of the target company.
5.2 Information Production
5.2.1 Redemption Rates
To test the first part of Hypothesis 1-1, we use the following regression:
where denotes the investment allocated to premium PIPE (at) investors as a fraction of the total IPO proceeds in SPAC , similarly for other allocations, and is a vector of covariates including total IPO proceeds and fixed effects that may vary across specifications. If premium investors generate value-relevant information, the higher the share allocated to premium investors, the lower the redemption rate, i.e., . For inference, we cluster the robust standard errors at the sponsor level to capture correlations within a sponsor but across investors and SPACs.
Table 4 provides evidence consistent with the hypothesis. In column (1), we include only the allocations to four types of PIPE investors. In column (2), we control for the size of the SPAC IPO and the liquidation risk of the SPAC when the target is announced using Days Remaining until the expiration. In column (3), we control for the quarter-fixed effects. In column (4), we control for the target company’s sector fixed effects to alleviate the problem that investors may have varying investment preferences across industries.
| (1) | (2) | (3) | (4) | (5) | |
| PIPE (pre) (%) | 0.125 | 0.192 | 0.148 | 0.156* | 0.148 |
| (0.123) | (0.119) | (0.116) | (0.095) | (0.094) | |
| Premium PIPE (at) (%) | -0.419*** | -0.380*** | -0.366*** | -0.214*** | -0.219*** |
| (0.069) | (0.064) | (0.059) | (0.050) | (0.051) | |
| Non-Premium PIPE (at) (%) | -0.013 | -0.009 | -0.002 | -0.037 | -0.036 |
| (0.034) | (0.035) | (0.034) | (0.030) | (0.031) | |
| PIPE (post) (%) | 0.120*** | 0.103*** | 0.114*** | 0.148*** | 0.148*** |
| (0.035) | (0.031) | (0.033) | (0.035) | (0.035) | |
| Days Remaining | ✗ | ✓ | ✓ | ✓ | ✓ |
| IPO Month FEs | ✗ | ✗ | ✓ | ✓ | ✓ |
| Merger Sector FEs | ✗ | ✗ | ✗ | ✓ | ✓ |
| Assortative Matching | ✗ | ✗ | ✗ | ✗ | ✓ |
| Adj. | 0.161 | 0.182 | 0.183 | 0.400 | 0.398 |
| Obs. | 444 | 444 | 444 | 430 | 430 |
Note: This table shows the results from estimating (1) using SPACs that went through an IPO from January 2010 to December 2021. The independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors.. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
We find that the allocations to premium PIPE (at) investors negatively correlate with the redemption rates, and this result is robust across all specifications. In particular, the estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE investors’ participation is associated with a 7.03 pps decrease in the redemption rate, which equals 12.4% of the sample mean.
A threat to identification is that the sample of sponsor and premium PIPE investors is not random. If better sponsors work with better investors and allocate more to those investors, the estimate of will suggest a larger (negative) decrease in the redemption rate. In other words, the estimate in column (4) may also capture the effect of matching where certain sponsors consistently secure favorable deals, which they then allocate to a select group of premium investors. To address this issue, we estimate the fixed effect of each sponsor and premium investor following the approach of Abowd, Kramarz and Margolis (1999) and include those fixed effects as additional controls in (1).
In particular, we estimate the following two-way fixed effect model of
that relates the money left on the table that investor gets from working with sponsor in time , , with their fixed-effects, and , respectively, and observed characteristics that includes quarter fixed effects, target sector fixed effects, the SPAC size, and the number of PIPE investors. Heuristically speaking, we use the returns, , to disentangle how much of it is attributable to the sponsor and investor. Most sponsors initiate more than one SPAC and work with different sets of investors each time. Likewise, investors work with different sponsors. These switching patterns identify these fixed effects.
A sponsor works with several premium investors, so to control for the quality of the sponsor and premium investors, for each SPAC, we use the sponsor fixed effect and the average fixed effect among all premium investors as additional regressors that capture the sponsor and premium investor qualities. Consistent with the matching hypothesis, these two fixed effects are positively correlated (regression coefficient , p-value ). In contrast, the sponsor fixed effect and the average fixed effect among all non-premium investors are negatively correlated (regression coefficient , p-value ).
The estimates with the additional controls are in Table 4 column (5). Comparing the estimates in columns (4) and (5), we can see that our result is qualitatively similar, i.e., premium investors provide certification roles. The estimates suggest that one standard deviation (32.85 pps) increase in premium PIPE (at) investors’ participation is associated with a 7.19 pps decrease in the redemption rate, which equals 12.7% of the sample mean.
We also find no association between non-premium PIPE (at) or PIPE (pre) allocations and redemption rates, as in Hypothesis 1-1 and Hypothesis 2-1. These results are consistent with the hypothesis that premium PIPE (at) investors are better informed about the deal quality than either the non-premium PIPE (at) investors or PIPE (pre).
The negative association between premium PIPE (at) allocation and redemption rates is consistent with premium investors producing information. Suppose not, and instead, suppose that sponsors underprice and distribute them to the premium investors in exchange for a promise to help them in the future on an unrelated SPAC. In that case, because the investors’ timing should be immaterial, we expect the allocations to PIPE (post) investors to correlate negatively with the redemption rate. However, as we see from Table 4, the two are positively correlated, further strengthening our finding.121212Relatedly, almost all PIPE (post) investors are non-premium investors. In fact, among the top 20 PIPE (post) investors by investment size, only Luxor Capital Group is a premium investor. Thus, it is reasonable to assume that sponsors include premium investors at the time of the announcement as certification and use non-premium investors to backfill the trust fund.
5.2.2 Announcement-day Return
To test the second part of Hypothesis 1-1 about the announcement-day return, we use the same model as (1) except the dependent variable is and the control variable is excluded. The estimation results are shown in Table 5. There is a strong positive association between announcement-day returns (%) and allocations to premium PIPE (at) investors. In particular, the estimates from our preferred specification in Table 5 column (5) suggest that one standard deviation (32.85 pps) increase in premium PIPE (at) investors’ participation is associated with a 4.14 pps increase in the SPAC’s announcement-day return, (110.7% of the sample mean). Thus, the effect of premium investors on announcement-day return is economically and statistically significant.
| (1) | (2) | (3) | (4) | (5) | |
| PIPE (pre) (%) | -0.118*** | -0.119*** | -0.099*** | -0.072 | -0.068 |
| (0.042) | (0.042) | (0.037) | (0.044) | (0.043) | |
| Premium PIPE (at) (%) | 0.148*** | 0.140*** | 0.136*** | 0.124*** | 0.126*** |
| (0.041) | (0.041) | (0.038) | (0.040) | (0.040) | |
| Non-Premium PIPE (at) (%) | -0.003 | -0.007 | -0.010 | -0.002 | -0.002 |
| Log IPO Proceeds | ✗ | ✓ | ✓ | ✓ | ✓ |
| Days Remaining | ✗ | ✓ | ✓ | ✓ | ✓ |
| IPO Month FEs | ✗ | ✗ | ✓ | ✓ | ✓ |
| Merger Sector FEs | ✗ | ✗ | ✗ | ✓ | ✓ |
| Assortative Matching | ✗ | ✗ | ✗ | ✗ | ✓ |
| Adj. | 0.099 | 0.104 | 0.098 | 0.095 | 0.091 |
| Obs. | 457 | 457 | 431 | 417 | 417 |
Note: This table shows the results from estimating (1), announcement-day return as the dependent variable, using SPACs that went through an IPO from January 2010 to December 2021. The key independent variable is the investment amount as a percentage of the SPAC’s IPO proceeds for all PIPE (pre) investors, premium PIPE (at) investors, non-premium PIPE (at) investors, and all PIPE (post) investors. Premium PIPE (at) investors are listed in Table 3. All other variables are defined in Table B1. All unbounded continuous variables are winsorized at their and percentile values. Robust standard errors are clustered at the sponsor level and are reported in parentheses. In column (5), we use 1,000 Bootstrap samples to estimate the standard errors. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
Furthermore, the announcement-day returns are uncorrelated with the allocations to non-premium PIPE (at) investors. Again, this finding is consistent with the hypothesis that larger investment by premium PIPE (at) investors acts as a certification of the quality of the deal. In contrast, investment by non-premium PIPE (at) is non-informative.
We also observe a negative correlation between announcement-day returns and PIPE (pre) investors’ allocations. Specifically, the estimates in column (5) suggest that one standard deviation (11.22 pps) increase in PIPE (pre) investors’ participation is associated with 0.84 pps lower in announcement-day return, which equals 22.5% of the sample mean.
Likewise, the hypothesis predicts a non-positive market reaction to the announcement of PIPE (post) investments. To assess this claim, we use an event study approach. In particular, we estimate the cumulative abnormal returns (CAR) (Campbell, Lo and MacKinlay, 1997) associated with announcements about PIPE investors while also announcing the target. The estimation period is from the SPAC’s unit separation date to the event window.
Table 6 reports the average of the CAR and their respective standard deviation across all SPACs using a three-day, five-day, and 10-day window around the event date. For each event window, we estimate three different asset-pricing models: CAPM, Fama-French three-factor, and Fama-French five-factor models. The average CAR estimates are not statistically different from zero in all cases. This result suggests that the investors do not respond to the announcement of PIPE (post), likely because almost all of them are non-premium and unlikely to add information.
| Event Window | CAPM | FF3 | FF5 |
| [-3, 3] | 0.71 | 0.74 | 0.75 |
| (0.80) | (0.79) | (0.78) | |
| [-5, 5] | -1.04 | -1.04 | -1.06 |
| (1.14) | (1.13) | (1.12) | |
| [-10, 10] | -2.37 | -2.30 | -2.28 |
| (1.92) | (1.92) | (1.91) |
Note: This table reports the SPAC share’s cumulative average abnormal return in percentage for three different event windows around announcements of 81 PIPE (post) transactions. Here, CAPM stands for capital asset pricing model, and FF3 and FF5 stand for Fama-French three- and five-factor models, respectively. The cross-sectional standard deviations of the cumulative abnormal returns are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
5.3 Quid Pro Quo
Next, we test Hypothesis 2 that there is a quid pro quo arrangement between sponsors and non-premium PIPE (at) investors. Recall that under favoritism, the current MLOT should be negatively correlated with the past MLOT. In particular, if the sponsor allocates securities to non-premium investors as a favor, and if the past MLOT is high, the sponsor will likely ask this investor to finance a weak deal today. We verify that the probability that a sponsor allocates securities to such a “relationship” investor increases if the investor has higher MLOT from the past deals and the SPAC deal is cold. Lastly, we show that quid pro quo lowers the liquidation risks.
5.3.1 Mean-reverting MLOT
As a first step, we assess if the MLOT exhibits mean reversion for the non-premium PIPE (at) investors, which is a testable implication of quid pro quo. In contrast, under the information production hypothesis, the correlation between the current MLOT and past MLOT should be zero or positive if there is an information spillover (Benveniste et al., 2003).
However, we need repeated deals between a sponsor and a PIPE investor to test the implication. In Table 7, we list the top 20 sponsors in our sample based on the total number of initialized and completed SPAC deals. In our sample, 139 sponsors have initialized more than one SPAC, totaling 439 SPACs. Furthermore, non-premium PIPE investors are more likely to participate in repeat deals (with the same sponsor) than premium investors.
To gauge the difference, consider the following. On average, the probability that a non-premium investor from the previous deal will participate in the current deal is 17%. There are around eight non-premium investors in a SPAC, so one of these eight non-premium investors in the current deal is from the previous one. If these investors were randomly (uniformly) participating in a SPAC deal, the probability would be much smaller because there are 1,792 non-premium investors in our sample. If non-premium investors participate randomly, the probability that at least one non-premium investor participated previously is negligible.
| Initialized | Completed | ||
| Sponsor | #SPAC | Sponsor | #SPAC |
| The Gores Group | 13 | The Gores Group | 9 |
| Cohen & Company | 11 | Chardan | 7 |
| Social Capital | 10 | Social Capital | 6 |
| Hennessy Capital Group | 8 | Cohen & Company | 6 |
| Cantor Fitzgerald | 8 | TPG | 5 |
| TPG | 8 | Cantor Fitzgerald | 5 |
| Michael Klein | 8 | Riverstone Investment Group | 5 |
| Jonathan J. Ledecky | 7 | Hennessy Capital Group | 5 |
| Chardan | 7 | Harry L. You | 5 |
| Hedosophia | 6 | Eagle Equity Partners | 5 |
| Harry L. You | 6 | Michael Klein | 4 |
| Riverstone Investment Group | 6 | Jonathan S. Huberman | 4 |
| Fortress Investment Group | 6 | Hedosophia | 4 |
| GigCapital Global | 6 | GigCapital Global | 4 |
| Apollo Global Management | 6 | Chinh Chu | 4 |
| Bill Foley | 6 | Niccolo de Masi | 4 |
| Chinh Chu | 5 | Bill Foley | 4 |
| Jaws Estates Capital | 5 | Apollo Global Management | 4 |
| Craig-Hallum Capital Group | 5 | Union Acquisition Group | 3 |
| Niccolo de Masi | 5 | Casdin Capital | 3 |
Note: This table presents the top 20 sponsors in terms of the total number of SPACs initialized and completed.
The probability gap reverses for the premium investors. The probability that a premium investor from the previous deal will participate in the current deal is 36.60%. There are, on average, three premium investors in a SPAC, so a sponsor includes one premium investor from the previous deal. There are 34 premium investors in our sample, so if they participate randomly in a SPAC, then the probability that at least one premium investor in the current deal has participated in the previous deal by the same sponsor is 90.31%.
Let be the average MLOT for all PIPE (at) investors in our sample. Table 8 is the summary statistics. On average, 28% of SPACs have at least one premium PIPE investor. Then, we estimate the following linear model,
where MLOT is the MLOT that PIPE (at) investor gets in the current SPAC deal initialized by the sponsor , and is the MLOT from the most recent deal in the past. As before, includes several covariates, including fixed effects (see Table 9).
| Mean | Std.Dev. | P10 | P50 | P90 | Obs. | |
| At Least 1 Premium | 0.28 | 0.45 | 0.00 | 0.00 | 1.00 | 5,812 |
| All PIPE Investors | ||||||
| Investment Amount ($mm) | 16.67 | 37.62 | 1.00 | 7.00 | 35.00 | 5,535 |
| MLOT ($mm) | 9.76 | 75.33 | -0.98 | 0.28 | 13.05 | 4,816 |
| No. of Past SPACs | 0.13 | 0.47 | 0.00 | 0.00 | 0.00 | 5,812 |
| MLOT (past) | 2.48 | 23.66 | 0.00 | 0.00 | 0.00 | 5,812 |
| Premium PIPE Investors | ||||||
| Investment Amount ($mm) | 20.50 | 38.79 | 3.00 | 10.00 | 45.00 | 1,609 |
| MLOT ($mm) | 11.01 | 45.60 | -1.17 | 0.58 | 23.97 | 1,462 |
| No. of Past SPACs | 0.20 | 0.53 | 0.00 | 0.00 | 1.00 | 1,647 |
| MLOT (past) | 3.98 | 19.82 | 0.00 | 0.00 | 2.55 | 1,647 |
| Non-premium PIPE Investors | ||||||
| Investment Amount ($mm) | 15.10 | 37.01 | 0.75 | 5.00 | 30.00 | 3,926 |
| MLOT ($mm) | 9.22 | 85.10 | -0.96 | 0.17 | 10.37 | 3,354 |
| No. of Past SPACs | 0.10 | 0.43 | 0.00 | 0.00 | 0.00 | 4,165 |
| MLOT(past) | 1.88 | 25.00 | 0.00 | 0.00 | 0.00 | 4,165 |
Under information production, the deviation of lagged MLOT from its mean is either uncorrelated or positively correlated with the current MLOT; hence in (3). Under quid pro quo arrangements, however, the MLOT should increase if the past deal was lower than the average and vice versa; hence, . The results are in column (1) of Table 9. As expected under favoritism, MLOT exhibits mean reversion.
| MLOT | |||
| (1) | (2) | (3) | |
| All PIPE Investors | -0.099* | ||
| Premium PIPE Investors | -0.025 | 0.235** | |
| (0.062) | (0.109) | ||
| Non-premium PIPE Investors | -0.218*** | -2.729*** | |
| (0.076) | (0.756) | ||
| Instrumental Variables | ✗ | ✗ | ✓ |
| F-stat (first-stage) | 92.82 | ||
| Adjusted | 0.418 | 0.419 | |
| Observations | 3,790 | 3,790 | 3,790 |
| Mean | Std.Dev. | ||
| (lag-MLOT-) | -8.12 | 5.41 | |
| (lag-MLOT-)(lag-MLOT-) | 121.96 | 85.54 | |
| (lag-MLOT-) | 135.78 | 5.77 | |
| (lag-MLOT-) | 166.93 | 297.54 | |
Note: This table shows results from estimating (4) using SPACs that went through IPO between January 2010 to December 2021. Each observation is at (SPAC, sponsor, PIPE investor) level. The dependent variable, MLOT, measures the money left on the table the PIPE investor makes from the SPAC deal. The key independent variables are the MLOT earned with the same sponsor in the last deal for all PIPE investors, premium PIPE investors, and non-premium PIPE investors. Additional controls include the log of IPO proceeds, the number of PIPE investors, the number of premium PIPE investors, days remaining, and separate fixed effects for the month, sponsor, and PIPE. All continuous variables are winsorized at the and percentile values. All other variables are defined in Table B1. Robust standard errors are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
Next, we differentiate between premium and non-premium investors and estimate
where is a binary variable equal to one if the PIPE (at) investor is a non-premium investor. The results from the estimation are in column (2) of Table 9. Comparing the second and third rows, we find that MLOT exhibits mean reversion for non-premium PIPE investors but not for premium investors. This result is consistent with the previous results that only the premium PIPE investors generate information, and Hypothesis 1-2.
A potential concern with the estimate of in column (2) of Table 9 is that sponsors’ decisions to include non-premium investors may be correlated with unobserved deal quality. Indeed, favoritism implies that the sponsor is more likely to involve non-premium investor(s) for a weak deal. If so, the sample used to identify the parameter is selected and not representative of the population. Because the sponsor is likely to allocate higher MLOT to non-premium investors in weak deals, the above estimate can be biased upwards.
We use the instrumental variable (IV) approach to address this problem. The idea behind our choice of IV is based on the following observation: suppose is endogenous because is endogenous, then can be an IV for under appropriate additional assumptions. For a recent application of this intuition, see Fabra and Reguant (2014) and Aryal, Ciliberto and Leyden (2022). In our setting, is and is the Non-Premium dummy, and given the specification in (4) is uncorrelated with the quality of the current deal.
However, may not satisfy the exclusion restriction because it directly affects the outcome variable. Based on the intuition above, we propose to use mean deviation from the second-to-last deal, i.e., as an excluded IV. Lagged variables are widely used as IVs in dynamic panel data settings under the Arellano-Bond estimator. Furthermore, to capture the nonlinearity of the interaction terms, we include as additional excluded variables.
The results from this estimation are in column (3) of Table 9. The first-stage Cragg-Donald Wald F-statistic of 92.82 suggests that our excluded variables are not weak instruments. Comparing these estimates with the ones in column (2), we see that, as expected, the fixed-effect estimate for the non-premium investor is biased upwards. More importantly, and consistent with the quid pro quo hypothesis, we find that one dollar higher mean deviation of MLOT in the last deal leads to a $2.729 reduction in the mean deviation from the current deal and vice versa presumably because now the non-premium investor is “asked” to participate in tepid deals that have smaller underpricing. In contrast, for the premium PIPE investors, the effect is the opposite: if the investor made one more dollar in MLOT above the mean in the last deal, on average, the MLOT in the next deal with the same sponsor increases by 23 cents consistent with the hypothesis that premium PIPE investors generate value-relevant information.131313Some deals have more than one sponsor. In such cases, we treated each sponsor separately. Although we do not show this here, we verify that the results are qualitatively the same if we exclude any SPAC with more than one sponsor or consider only the sponsor with the maximum number of past SPAC deals.
5.3.2 Allocations and Past Relationship
Next, we test the Hypothesis 2-3. A sponsor builds a relationship with non-premium investors by providing MLOT from good deals, so if and when the new target firm is weak, it expects the investors with high past MLOT, i.e., the relationship investors, to return the favor by investing in the current deal. To this end, we evaluate if a sponsor is more likely to use non-premium investors if their past MLOT from the sponsor is high and the current target is weak.
Although we do not observe the quality of the target, we propose to proxy it by the difference between the SPAC’s offer share price, i.e., the share price at the SPAC IPO, and the share price on the closing day of the merger. The larger the gap, the lower the quality.
Then, we estimate the probability that a sponsor allocates securities to a non-premium investor as a function of the past MLOT and a measure of cold, among other variables. For a sponsor , let denote the set of relationship investors as the set of all non-premium investors that invested in at least one SPAC with sponsor . Let be a binary variable equal to one if investor is allocated securities from SPAC started by sponsor , and zero otherwise. Then we estimate the participation probability, , under the assumption that it takes a logit form, with
where is the vector of controls shown in Table 9; is the log of the total MLOT made in past deals with and measures the past relationship between and ; and, and as we defined above, is the measure of the quality of the deal .
We estimate (5) using the method of maximum likelihood and present the results in column (1) of Table 10. The estimated coefficient of the interaction term (third row) in column (1) is positive (p-value ), supporting the quid pro quo hypothesis that the sponsor first builds a relationship and draws it down as and when needed.
| Participation | Allocation | ||||
| (1) | (2) | (3) | (4) | (5) | |
| 0.085 | -0.004 | -0.088 | -0.335 | 0.019 | |
| (0.136) | (0.123) | (0.098) | (0.081) | (0.010) | |
| [0.013] | |||||
| Cold | 0.006 | 0.004 | -0.003 | 0.003 | -0.00003 |
| (0.001) | (0.001) | (0.002) | (0.002) | (.00006) | |
| [0.0006] | |||||
| 0.008 | 0.011 | 0.011 | 0.008 | 0.0003 | |
| (0.002) | (0.003) | (0.003) | (0.002) | (0.0001) | |
| [0.00012] | |||||
| Outside Option | -0.017 | -0.007 | -0.002 | ||
| (0.003) | (0.002) | (0.002) | |||
| 0.239 | |||||
| [0.024] | |||||
| Log IPO proceeds | ✗ | ✓ | ✓ | ✓ | ✓ |
| Days Remaining | ✗ | ✓ | ✓ | ✓ | ✓ |
| No. Premium Investors | ✗ | ✓ | ✓ | ✓ | ✗ |
| Sponsor Dummies | ✗ | ✗ | ✓ | ✓ | ✗ |
| Month Dummies | ✗ | ✗ | ✗ | ✓ | ✓ |
| Pseudo/Adj. | 0.0098 | 0.0644 | 0.2763 | 0.4061 | 0.8114 |
| Obs. | 10,104 | 10,044 | 7,917 | 7,913 | 7,838 |
Note: Columns (1) to (4) of this table show the MLE of participation (5), and column (5) shows the estimates of the allocation (6). Each observation is at (SPAC, sponsor, non-premium PIPE investor) level. For each SPAC, non-premium investors include all relationship investors who participate in at least one SPAC deal initialized by the sponsor during the sample period. For the first four columns, the dependent variable, , equals one if the non-premium investor participates in the focal SPAC deal. The dependent variable for the last column, , is the investor’s allocation as a fraction of all relationship investors’ allocations. The key independent variables include the non-premium investor’s MLOT earned with this sponsor in the past in log form, , and a measure of the (reciprocal) of the deal quality, Cold. For the allocation regression, we also include the estimated probability of participation using estimates from column (4) as an additional variable. All continuous variables at winsorized at percentile. Robust standard errors are clustered at the sponsor-PIPE pair level and reported in parentheses. For column (5), Bootstrap standard errors (500 replications) are reported in the square brackets. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
One possible concern with the estimation is that it ignores the outside option of the sponsor, which may be correlated with the unobserved error. A better outside option lowers the probability of issuing securities to a non-premium investor and vice versa. Ignoring outside options could introduce attenuation bias if cold deals have profitable outside options. To alleviate this concern, we use the highest MLOT among all premium PIPE investors not included in the current deal as an additional control variable to capture the strength of the outside option. We also include the log of IPO proceeds, days remaining, and the number of premium investors as additional covariates.
The estimates from this new model are in column (2) of Table 10. A better outside option lowers the probability of selecting the non-premium investor, and our previous finding becomes slightly stronger. The results are robust when we include dummy variables for the sponsor and the month when the deal is announced. In all cases, the coefficient for the interaction term is positive and estimated precisely (p-value ).
To help interpret these estimates, we evaluate the probability in (5) using the estimates from column (4) of Table 10 at different levels of past MLOT and quality of the target, and hold other variables at their average value. Investors with a stronger relationship have higher participation rates in deals with lower demand. In SPAC deals with Cold at its -percentile, non-premium investors who have a strong relationship with the sponsor (one standard deviation above the mean) are 3.54 pps more likely to participate than in SPAC deals with Cold at its -percentile. Conversely, non-premium investors without a prior relationship with the sponsor are less likely to participate in cold deals.
Next, we examine if, conditional on participation, the sponsor allocates a higher portion of securities to non-premium investors with a stronger relationship when a deal is cold and vice versa. To this end, we determine the allocation of SPAC ’s securities from sponsor to PIPE investor among all relationship investors , , as the ratio of ’s shares to the total shares allocated to . We set the share to zero if a non-premium investor does not participate in a deal. The mean allocation is , and the standard deviation is . Furthermore, the allocation is skewed towards zero, with the median of and the -percentile at . Given that there are 1,792 non-premium investors, and so the allocation in our sample is skewed to the left with a mass at zero, we estimate the following Tobit model:
where is always observed but the latent variable is observed only when . We further assume that is independent of and that is a mean-zero Normal random variable with unknown variabce and . Furthermore, is the estimated participation probability of allocation given and is determined by the estimates in column (4) of Table 10. Thus, the outside option serves as our excluded variable, with the underlying exclusion restriction being that the outside option should not directly affect a non-premium investor’s allocation.
In Column (5) of Table 10, the estimates indicate a positive association between a higher probability of participation and an increased allocation of shares. Moreover, non-premium investors who have previously profited from deals with the sponsor receive a larger share of allocation than their counterparts when a deal is considered cold.
In economic terms, a one standard deviation increase in the interaction term leads to a 5.8% increase in allocation relative to its standard deviation, amounting to 14.80% of the sample mean. In our sample, an average non-premium investor is allocated 2.68% of the total non-premium investors’ allocations. Furthermore, as the non-premium investor’s MLOT (past) times the deal’s Cold measure increase together by one standard deviation, the allocation fraction rises by 0.40 pps to 3.08%. In terms of dollar amount, the allocation increases from $0.40 million to $0.46 million. These findings support the hypothesis that non-premium investors and sponsors engage in quid pro quo agreements.
5.3.3 Relationship Capital and Liquidation Risk
We have shown that the relationship investors support the sponsor by participating in a weak deal and purchasing a larger fraction of shares. These non-premium investors received larger MLOTs from the same sponsor in the previous deals. Therefore, it stands to reason that a sponsor builds these relationships with an eye toward weak deals in the future. A sponsor with more relationship capital can afford to take more risk by working with a weak target firm because it can draw down the relationship capital.
Next, we test Hypothesis 2-4 by estimating the liquidation probability as a function of MLOT earned by premium and non-premium investors in previous deals initiated by the same sponsor using a linear probability model. In our sample, we observe 191 liquidations and 473 successful mergers, and we want to test if the probability of liquidation decreases with the sponsor’s “relationship payments” to non-premium investors.
We regress the binary variable that is equal to one if the SPAC is liquidated and zero otherwise, on the past MLOT, the premium and non-premium investors have made from the current sponsor and several other controls.
| (1) | (2) | (3) | (4) | (5) | |
| Premium Investors | 0.057** | 0.049 | 0.054* | 0.059* | 0.060* |
| (0.029) | (0.033) | (0.032) | (0.033) | (0.033) | |
| Non-premium Investors | -0.064** | -0.075** | -0.077** | -0.079*** | -0.081*** |
| (0.029) | (0.030) | (0.030) | (0.030) | (0.031) | |
| Controls | ✓ | ✓ | ✓ | ✓ | ✓ |
| IPO Month FEs | ✗ | ✓ | ✓ | ✓ | ✓ |
| Target Sector FEs | ✗ | ✗ | ✓ | ✓ | ✓ |
| Jurisdiction FEs | ✗ | ✗ | ✗ | ✓ | ✓ |
| Sponsor Quality | ✗ | ✗ | ✗ | ✗ | ✓ |
| Adj. | 0.047 | 0.121 | 0.127 | 0.126 | 0.123 |
| Obs. | 806 | 781 | 758 | 757 | 757 |
Note: This table demonstrates that SPACs initiated by sponsors with a history of stronger relationships with non-premium investors are less likely to liquidate when a target is announced. All regressions use linear probability models. Dependent variable is a dummy variable that is one if the SPAC liquidates and zero otherwise. The primary independent variables are the total MLOT earned by premium and non-premium investors in previous SPACs initiated by the same sponsor. Controls include (log) IPO proceeds, number of past successful deals, dummy variable if there were PIPE (pre) investors, and number of active SPACs. Sponsor quality is the sponsor’s fixed effect estimated from (2). All continuous variables are winsorized at the and percentiles. Robust standard errors, clustered at the sponsor level, are reported in parentheses. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
The results from this estimation exercise are in Table 11. Column (1) suggests that SPACs initialized by sponsors who have built a stronger relationship with non-premium investors are less likely to liquidate. We progressively introduce additional control variables from columns (2) to (5). In column (2), we include SPAC’s IPO month fixed effect to capture time differences. In column (3), we include fixed effects for the SPAC’s targeted sector at its IPO to capture systematic differences across sectors that may affect the liquidation risk. In column (4), we incorporate SPAC’s jurisdiction fixed effects based on the jurisdiction the SPAC is registered in to capture differences in the legal systems that may influence the liquidation rules. Finally, in column (5), we included the sponsor fixed effect we estimated from Equation (2) as a proxy for the sponsor’s quality.
Column (5), our preferred specification, suggests that doubling the sponsor’s relationship payments to non-premium investors decreases liquidation probability by 8.1%, which is 33.6% of the sample average liquidation rate, and this estimated effect is statistically significant at the 1% confidence level.141414 Although not reported, we verify that even when we use the competing-risk model to account for right censoring because 16% of the announced deals in our sample are still searching for a target, our finding holds, albeit the estimates are slightly nosier than in Table 11.
In summary, we find evidence consistent with these conditions: (i) premium PIPE investors generate value-relevant information, (ii) non-premium PIPE investors are paid agency fees, and (iii) the non-premium investors help close weaker deals in the future. In Appendix A, we show that these results are robust to removing outlier investors, placebo treatments, and the definition of premium PIPE investors.
6 Observed and Unobserved Heterogeneity
In this section, we determine how much of the difference in average earnings between premium and non-premium investors can be explained by differences in their observable characteristics. If they have different “business models” (information production vs. agency service), we expect observed heterogeneity to explain a smaller portion of the difference.
To minimize the effect of outliers in this exercise, we exclude seven out of 473 successful SPACs with closing share prices higher than $100.151515These SPACs are: Queen’s Gambit Growth Capital, Insurance Acquisition Corp., Matlin & Partners Acquisition Corporation, Northern Genesis Acquisition Corp. II, VPC Impact Acquisition Holdings III, Inc., Reinvent Technology Partners Z (f/k/a Reinvent Technology Partners B), and Ajax I. In Figure 2 (left panel), we display the histogram of total MLOT paid to all PIPE (at) investors across SPACs that have successfully merged with a private target firm. The average MLOT per deal is $62.38 million, with a large standard deviation of $199.07 million. These earnings are divided among, on average, 10.76 PIPE investors (see Table 2), of which 3.02 are premium investors. In Figure 2 (right panel), we show the shares of premium PIPE investors (left vertical line) and the total MLOT (right vertical line). Averaging these shares, we find that premium PIPE investors get (around) 26% of the MLOT. In other words, 26% of the underpricing at the time of the business closing announcement is for information production, and 74% is for quid pro quo.
Note: The left panel is the histogram of total MLOT for all investors. The right panel shows the premium PIPE investors’ share of MLOT (vertical axis on the left) and the total MLOT in that deal (vertical axis on the right).
Considering each investor separately, the MLOT of a premium investor, on average, is $6.96 million, and that of a non-premium investor is $3.17 million. Next, we determine what part of this $3.79 million difference in average MLOT can be explained by differences in the observable characteristics of these two types of investors.
First, we introduce new variables capturing some of the observed differences between these two groups. In particular, for each PIPE investor and each SPAC, we determine the number of SPACs they participated in before the current one, how much they invested, the number of past SPACs that involved a sponsor different from the current one, total MLOT from other sponsors, the number of SPACs with sponsors different from the current one, and whether the target is in the same industry as the current target. These variables capture market-wide experience and successes.
We have 4,152 deals at the SPAC and PIPE (at) investor level, of which 1,245 are with premium investors and the rest with non-premium investors. Table 12 presents the descriptive statistics and the results from the tests of mean differences between premium and non-premium investors.
Premium PIPE Sample |
Non-Premium PIPE Sample |
Difference |
MLOT ($mm) |
6.95 |
3.17 |
3.78 |
| SPAC Characteristics | |||
IPO Proceeds ($mm) |
397.25 |
363.87 |
33.38 |
Days Remaining |
492.98 |
467.59 |
25.38 |
| PIPE Investor Characteristics | |||
No. of past SPACs |
23.99 |
5.51 |
18.48 |
Past Investment Amount ($mm) |
460.09 |
50.77 |
409.32 |
No. of SPACs with other sponsors |
24.40 |
5.44 |
18.95 |
Past MLOT with other sponsors ($mm) |
292.85 |
24.61 |
268.25 |
No. of SPACs with other sponsors, same industry |
3.59 |
0.80 |
2.79 |
Past MLOT with other sponsors, same industry ($mm) |
45.83 |
3.65 |
42.18 |
Note: This table reports means and differences in our sample’s average characteristics for premium and non-premium investors. , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
The two groups display distinct MLOTs and characteristics. Specifically, premium investors are involved in larger SPACs requiring fewer days to identify a target. Moreover, these investors possess greater experience in the SPAC market, as evidenced by a higher number of previous SPAC participations (23.99 vs. 5.51) and a larger investment amount ($460.09 million vs. $50.77 million). They also work with more SPACs (24.40 vs. 5.44) and obtain a higher MLOT ($292.85 million vs. $24.61 million) from other sponsors. In addition, premium investors operate more in the same industry with different sponsors (3.59 vs. 0.80).
Next, we use the Kitagawa-Blinder-Oaxaca decomposition to document the extent and drivers of the MLOT differences between premium and non-premium investors. In particular, we begin with the following: where is the vector of observed characteristics in Table 12, and is a mean zero random variable. Let , where denotes the sample average of the group . We can then re-write as
The explained portion is the portion of the MLOT difference explained by differences in observed characteristics between premium and non-premium investors. In other words, if we change to , it measures how much additional MLOT a non-premium investor gets. The unexplained portion is the part driven by differences in the coefficients.
Note: The figure displays the explained components of a twofold Kitagawa-Blinder-Oaxaca decomposition of the premium vs. non-premium MLOT. The x-axis is the mean difference in MLOT (in millions), and the y-axis lists observable characteristics.
We find that the mean increase in non-premium investors’ MLOT is $0.04 million (s.e. $1.38 million) if they had the same characteristics as premium investors. So, the differences in observed characteristics do not account for the MLOT gap. Figure 3 displays how much each of the eight components in explains, along with their confidence intervals. We find that the difference in average MLOT cannot be explained by differences in average , suggesting these two types of investors have different business models.
7 Conclusion
Understanding reasons for underpricing hinges on the relative importance of information production and quid pro quo. In this paper, we determine the relative importance within the context of SPACs. Distinguishing between premium and non-premium PIPE (institutional) investors, we show that the former produce information and, on average, are paid 26% of the total underpricing. The latter extracts the remainder 74% as quid pro quo payments.
While we have relied on some institutional features specific to SPACs, understanding how these features arise to alleviate information and agency frictions has implications for security issuance under asymmetric information in general and IPOs in particular. After all, the SPAC business combination process is also a case of security issuance through sponsors as intermediaries, with PIPE investors as buyers and the target company as the seller.
There are two main takeaways for IPOs. First, information production and quid pro quo can play important roles in IPO underpricing. In other words, they are not necessarily mutually exclusive. Second, non-premium PIPE investors help sponsors with weak deals, suggesting that, on average, they are value-enhancing for targets and issuers. This role is unlike the traditional IPO, where some investment bank clients get favorable allocations of “hot” IPO shares in exchange for paying high trading fees to the bank.
The benefit of quid pro quo payment is a function of its positive effect on successful business combinations. However, to determine if quid pro quo payments enhance welfare, we have to show that the total welfare is higher when the sponsor is allowed to build and deplete relationships than if not. These are interesting questions for future research.
APPENDIX
Appendix A Robustness
In Section 5, we found evidence consistent with the hypotheses that premium investors produce value-relevant information and non-premium investors engage in quid pro quo arrangements. We also found evidence that the quid pro quo relationship between a sponsor and non-premium investors enables weaker private firms to go public. In this section, we perform a series of robustness exercises.
A.1 Possible Outliers.
Note: The figure shows the scatter plot of SPACs by month and year. The top figure shows the number of SPAC mergers, and the bottom figure shows the deal sizes (total IPO proceeds) in billions of dollars. The size of each dot is proportional to the number of deals in a month (top) and the total IPO proceeds (bottom) in that month.
As shown in Figure A4, although our sample starts in 2010, most variation comes from SPAC from 2020 to 2022. Although our sample includes the so-called boom and bust cycle for SPAC, we may worry about the effects of outliers in our analysis. One such concern is that not all 34 premium PIPE investors (in Table 3) produce value-relevant information and that our results are driven by either the large premium PIPE investors or that the small premium PIPE investors are, in fact, non-premium investors.
To address this concern, we exclude six (large) premium investors with at least $1 billion invested in SPACs and re-estimate all our models. Then, we repeat the same exercise excluding seven (small) premium investors with at most $400 million invested in SPACs.
| Redemption | Announcement-day | Mean Reversion | Liquidation | |||||
| Rate (%) | Return (%) | MLOT | ||||||
| (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) | |
| PIPE (pre) | 0.174* | 0.151 | -0.082* | -0.070 | ||||
| (0.098) | (0.093) | (0.044) | (0.043) | |||||
| Premium PIPE (at) | -0.244*** | -0.219*** | 0.144** | 0.126*** | 0.153 | 0.149 | 0.066** | 0.059* |
| (0.082) | (0.052) | (0.059) | (0.041) | (0.104) | (0.099) | (0.033) | (0.034) | |
| Non-premium PIPE (at) | -0.041 | -0.038 | 0.001 | -0.001 | -1.633*** | -2.075*** | -0.083*** | -0.078** |
| (0.031) | (0.031) | (0.014) | (0.015) | (0.531) | (0.615) | (0.029) | (0.031) | |
| PIPE (post) | 0.152*** | 0.148*** | ||||||
| (0.036) | (0.035) | |||||||
| Exclude Large Premium Investors | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ |
| Exclude Small Premium Investors | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ | ✗ | ✓ |
| Pseudo/Adj. | 0.386 | 0.397 | 0.069 | 0.090 | 0.127 | 0.125 | ||
| F-stat (first-stage) | 175.54 | 140.73 | ||||||
| Observations | 430 | 430 | 417 | 417 | 3,487 | 3,578 | 758 | 757 |
Note: This table presents estimates after excluding outliers in premium investors. Premium PIPE (at) investors are listed in Table 3. They include Fidelity Management & Research Company, BlackRock, Inc., Capital Research and Management Company, Alyeska Investment Group L.P., Morgan Stanley, and Millennium Management, LLC. Small premium investors are those that invested less than $0.4 billion and participated in at most 50 SPAC deals. They include Ghisallo Capital Management LLC, Jane Street Group, LLC, Park West Asset Management LLC, Kepos Capital LP, Linden Advisors LP, Schonfeld Strategic Advisors LLC, and BlueCrest Capital Management Ltd. Columns (1) and (2) have the same specification as Table 4, column (5). Columns (3) and (4) have the same specification as Table 5, column (5). Columns (5) and (6) have the same specification as Table 9, column (3). Columns (7) and (8) have the same specification as Table 11, column (5). , and denote p-values less than 0.1, 0.05, and 0.01, respectively.
The results are presented in Table A1. In particular, the estimates from excluding large premium investors are in columns (1), (3), (5), and (7) of Table A1, and the estimates from excluding small premium investors are in the remaining columns. Comparing these estimates with those in columns (5), (5), (3), and (5) of Tables 4, 5, 9, and 11, respectively, we see that our main findings are neither due only to the large premium investors nor the small premium investors.
A.2 Placebo Exercises
Note: This figure shows coefficient estimates for premium and non-premium PIPE investors, using randomly generated premium PIPE identities. We randomly pick 34 PIPE investors out of 1,826, label them as premium PIPE investors, and re-estimate our models. We repeat this exercise to have 500 sets of randomly chosen premium PIPE samples. In these five models, the specifications are the same as in Tables 4 and 5, column (5), Table 9 column (3), and Table 11 columns (5), respectively.
We use a placebo exercise to assess whether our choice of premium and non-premium investors is meaningful. To this end, we explore the effect of choosing an arbitrary set of PIPE investors and designating them as premium PIPE investors on the estimates. If our results were only noise, then a random group of premium PIPE investors should also give similar estimates. To implement this “placebo” exercise, we randomly picked 34 out of all 1,826 PIPE investors and labeled them premium investors. Then, we estimate all the models using this new set of premium investors. We repeat this exercise 500 times and present only the coefficients for premium and non-premium investors and the 95% confidence intervals in Figure A5, in the first and second columns, respectively. The five rows are redemption rate, announcement day return, mean-reversion, liquidation risk, and probability of liquidation.161616For the mean-reversion exercise, we present the first 500 cases with the first-stage F-stat .
For the redemption rates, comparing the first row in Figure A5 with the estimates in Table 4 column (5), we see that the randomly chosen premium PIPE investors do not affect redemption rates. However, because the non-premium PIPE investors include some (actual) premium PIPE investors, the redemption rate is low but negatively correlated with allocations for non-premium PIPE investors. Similarly, comparing the second row in Figure A5 with the estimates in Table 5 column (5), we find no effect of premium PIPE investors on the announcement-day return but a small and positive effect of non-premium PIPE investors on return because the latter set now contains the premium investors, see Table 3. These effects are consistent with only the premium investors in Table 3 producing information.
For the mean-reversion (in the third row), some estimates are negative, consistent with the new premium PIPE investor containing non-premium PIPE investors and vice versa. Finally, for the probability of liquidation, comparing rows four in Figure A5 with Table 11 columns (5), we see the estimates for non-premium investors are smaller in magnitude and not statistically significant.
Variable |
Unit |
Definition |
| Panel A: SPAC-Level Characteristics | ||
| SPAC Measures | ||
IPO proceeds |
MM USD |
Gross proceeds raised in the IPO, including any full or partial exercise of the Greenshoe. |
No. of Warrant |
Number of warrants included in the unit issued at the SPAC’s IPO. |
No. of Right |
Number of rights included in the unit issued at the SPAC’s IPO. |
Overallotment |
Amount of the overallotment option exercised by the underwriter(s) as a fraction of total IPO proceeds. |
Listing-day return |
SPAC IPO investors’ first-day return. |
Days searching |
Day |
The number of days between the SPAC’s IPO date and the target announcement date. |
Days remaining |
Day |
The number of days between the SPAC’s target announcement date and the liquidation deadline. |
Announcement-day return |
One-day return on the SPAC’s shares on the target’s announcement date. |
Redemption |
Redeemed SPAC common shares as a percentage of total shares issued at IPO. |
Days deSPAC |
Day |
The number of days between the SPAC’s target announcement date and the merger closing date. |
Cold |
pps |
The difference between the SPAC’s offer price and the SPAC’s share price on the closing day of the merger. |
Liquidation |
binary |
One if the SPAC liquidated during our sample period, zero otherwise. |
Target sector |
Sector which the SPAC’s registration statement indicates it will focus on in its search for a target for a business combination. |
Merger sector |
Sector in which the merger company operates. |
Jurisdiction |
State (if incorporated in the U.S.) or country in which the SPAC is incorporated at the time of the IPO. |
No. active SPACs |
Number of SPACs actively searching for targets at the time of the focal SPAC’s IPO. |
No. past successful deals |
Number of SPACs that completed the merger process before this focal SPAC’s IPO, and initiated by the same sponsor. |
| PIPE Participation | ||
{PIPE (pre)} |
Binary |
One if PIPE (pre) investors participate in the deal. |
{PIPE (at)} |
Binary |
One if PIPE (at) investors participate in the deal. |
{PIPE (post)} |
Binary |
One if PIPE (post) investors participate in the deal. |
{Premium PIPE (at)} |
Binary |
One if premium PIPE (at) investors participate in the deal. |
{Non-Premium PIPE (at)} |
Binary |
One if non-premium PIPE (at) investors participate in the deal. |
No. of PIPE (at) |
Number of PIPE (at) investors. |
No. Premium PIPE (at) |
Number of premium PIPE (at) investors. |
No. of Non-Premium PIPE (at) |
Number of non-premium PIPE (at) investors. |
| PIPE Investment/IPO Proceeds | ||
%PIPE (pre) |
PIPE (pre) investors’ investment as a percentage of total IPO proceeds. |
%Premium PIPE (at) |
Premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
%Non-Premium PIPE (at) |
Non-premium PIPE (at) investors’ investment as a percentage of total IPO proceeds. |
%PIPE (post) |
PIPE (post) investors’ investment as a percentage of total IPO proceeds. |
| Panel B: PIPE-Investor-Level Characteristics | ||
Non-Premium |
Binary |
One if either the PIPE (at) investor’s number of participated PIPE deals or the total commitment amount is not in the top 5th percentile. |
Investment Amount |
MM USD |
Dollar amount of the PIPE investor’s investment in the SPAC deal. |
MLOT |
MM USD |
The difference between the value of securities held by the PIPE investor on the first day of listing of the post-merger company and the PIPE investor’s initial investment. |
lag-MLOT |
MM USD |
The PIPE investor’s MLOT from the previous SPAC deal with the same sponsor. |
lag-MLOT |
MM USD |
The PIPE investor’s MLOT from the one before the previous SPAC deal with the same sponsor. |
MLOT(past) |
MM USD |
Total MLOT a PIPE investor made in past deals with the sponsor. |
No. of past SPACs |
Number of SPAC deals participated by the PIPE investor before the focal SPAC deal. |
Past Investment Amount |
MM USD |
The PIPE investor’s total amount of investments in SPAC deals before the focal SPAC deal. |
No. of SPACs with other sponsors |
Number of other sponsors’ SPAC deals participated in by the PIPE investor before the focal SPAC deal. |
Past MLOT with other sponsors |
MM USD |
The PIPE investor’s total MLOT made in other sponsors’ SPACs before the focal SPAC deal. |
No. of SPACs with other sponsors, same industry |
Number of other sponsors’ SPAC deals that found targets in the same industry as the focal SPAC and participated by the PIPE investor before the focal SPAC deal. |
Past MLOT with other sponsors, same industry |
MM USD |
The PIPE investor’s total MLOT made in other sponsors’ SPACs that found targets in the same industry as the focal SPAC. |
Outside Option |
MM USD |
Highest MLOT among all premium PIPE investors not included in the current deal. |
Participation |
Binary |
One if the sponsor allocates securities to the non-premium PIPE investor, and zero otherwise. |
Allocation |
Ratio of the non-premium investor’s shares to the total shares allocated to the sponsor’s relationship investors. |